Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Laser Bench setup)
gosselin, derossi - 10:59 Tuesday 17 February 2026 (68699) Print this report
Comment to Slave laser temperture controller failure (68693)

Yesterday evening, for an unkown reason, the current of the pumping diodes suddently decreased (figure 1). 
This morning we went in EER, we had 27.8 A for diode 1 and 27.6 for diode 2. We brought the current back to 29 A and got the system as it was on saturday (figure 2). 
We waited for everything to thermalized and we went to realign the PMC, shift and tilt (figure 3) 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×